Cassini Significant Events 03/14/07 - 03/20/07
March 22, 2007
(Source: Cassini Project)
The most recent spacecraft telemetry was acquired on Tuesday, March 20, from the Goldstone tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" web page at http://saturn.jpl.nasa.gov/operations/present-position.cfm .
Wednesday, March 14 (DOY 073):
Science activities this week included the start of a Magnetospheric and Plasma Science campaign with observations of the afternoon magnetospheric boundaries at 35 Rs apoapsis distance and a range of inclinations. The Visual and Infrared Mapping Spectrometer led an Optical Remote Sensing F-ring movie which executed over 17 hours, the Magnetometer Subsystem performed one of its periodic calibration rolls necessary due to the instrument degradation, and the Composite Infrared Spectrometer (CIRS) performed temperature mapping and ring system photometry. Beginning on Monday, March 19, the Rings segment in Orbit 41 included latitude-phase coverage of the rings, and ring stellar occultations.
Instrument, Navigation, and Spacecraft Operations teams delivered their official input files today as part of the S31 Science Operations Plan update process.
A presentation on Titan and Enceladus science goals for the Cassini extended mission was presented to the flight team today. This was the second of four science presentations that were given as part of the Senior Review to NASA personnel back in February.
Thursday, March 15 (DOY 074):
The Encounter Strategy Meeting for Titan 27-28 was held today. The meeting covered the period of time from March 26 through April 10, and maneuvers 101-103.
The Cassini "Planetariums & More" web page has been updated. A new "Create your own Cassini Presentation" feature links multimedia,
downloads, education and other resources all on one page. Another
new feature - Cassini Museum Events in your area - lists museum
events by state. Finally, the 35-minute DVD Ring World 2 is still
available and is also online as a podcast in both English and Spanish. Here's the link:
Saturday, March 17 (DOY 076):
Orbit Trim Maneuver (OTM) #99 was performed today. This is the apoapsis maneuver setting up for the Titan 27 encounter on March 26.
The main engine burn began at 6:59 PM PST. Telemetry immediately after the maneuver showed the burn duration was 9.95 seconds, giving a delta-V of 1.60 m/s. All subsystems reported nominal performance after the maneuver.
Monday, March 19 (DOY 078):
The S28 DOY 084-085 Live Moveable Block kickoff meeting was held today. This update is being requested by Radio Science (RSS) and is based on the release of the OTM-99 solution from March 18. RSS has received the configuration files with new orbit determination and the update is currently a "GO". RSS cannot make a final Go/No-Go decision until Wednesday when it is decided if OTM #100 will be required or not.
The science operations plan update process for S32 kicked off today.
Tuesday, March 20 (DOY 079):
The S29 final sequence approval and command approval meetings were held today. All 13 instrument expanded blocks were uplinked to the spacecraft yesterday, the background sequence will be uplinked on Friday, March 23, and execution will begin on Wednesday, March 28.
The topic this week at the Mission Planning Forum was an update on dust hazards for prime mission, and introduction to dust hazard profile and options for extended mission.